How much heat is required to convert 1g of ice at 0 degree Celsius to steam at 100 degree Celsius (Heat of fusion of ice = 335 x 10^3 Jkg^-1, heat of vaporization of water = 2.26 x 10^6JKg^-1, specific heat capacity of water = 4200Jkg^-1K^-1)
1
Expert's answer
2019-05-27T11:15:01-0400

The total heat consists of the heat of ice melting, water heating and vaporization:

"Q=Q_m+Q_h+Q_v"

the heat of ice melting:


"Q_m=\lambda m"

Heat of water heating:


"Q_h=Cm(T_2-T_1)"

And heat of vaporization:


"Q_v=Lm"

Then


"Q=(\lambda+L+C(T_2-T_1))m=\\
=(335\times10^3+2.26\times10^6+4200\times100)\times0.001 = 3015J"